How to prove log2(7 + 1/n) = Big Omega(1)

I know that by definition to be Big omega of 1 the f(n) is not bounded by n to infinity, but its bounded to a constant C.
Answer #1
Big Omega Definition:
----------------------------
f(n)=Ω(g(n)) usually defined by saying for some constant c>0 and all large enough n, f(n)≥c g(n).


log2(7 + 1/n) = Big Omega(1)
Here, f(n) = log2(7 + 1/n)
For all value of n > 0 the condition log2(7 + 1/n) >= 2.8
So, We can say that log2(7 + 1/n) >= 2.8*1
Where c = 2.8 and n > 0 and g(n) = 1
So, f(n)=Ω(g(n))
f(n)=Ω(1)